One number is 5 more than another. five times the smaller equals 4 times the larger. Find the numbers.

The Colts played 76 games. They one 3 games as many games as they lost. How many did they win?

Answer by rapaljer(4671)

For your second problem, did you mean that they "won three times as many games as they lost"?? If so, then let x = number of games they lost.

Let x = number of games lost
3x = number of games won

x+ 3x = 76
4x = 76
x= 19 games lost
3x = 3(19) = 57 games won

It checks since 19 + 57 = 76 total games.

R^2 at SCC